# Pareto Set Learning for Multi-Objective Reinforcement Learning

### v1

@@ Introduction
Reinforcement Learning (RL) (Sutton and Barto 2018) has become a prevalent technique for addressing decision-making and control problems in a wide range of scenarios, including video games (Mnih et al. 2015), navigation (Yang et al. 2019), robotics (Schulman et al. 2017), dynamic algorithm configuration (Xue et al. 2022), and financial cloud services (Yang et al. 2024). In standard RL, the main objective is obtaining an optimal policy that achieves the best expected returns. However, in many real-world scenarios, spanning from robotics, where the balance between speed and energy efficiency is paramount (Todorov et al. 2012), to smart systems design, where the trade-offs between competing metrics must be dynamically managed (Azzouz et al. 2018), the problems may involve multiple conflicting objectives. Contrary to standard RL, Multi-Objective RL (MORL) aims to generate an optimal policy tailored to each preference over multiple objectives, constructing a diverse set of optimal policies that collectively form a Pareto set.
@@
Traditional MORL methods transform the multi-dimensional objective space into a scalar representation (Mannor and Shimkin 2001) through the static assignment of weights to each objective, achieving a trade-off between multiple objectives. Such approaches, however, struggle with objectives of varying magnitudes and require extensive domain knowledge to set the appropriate weights/preferences (Amodei et al. 2016). Moreover, they inherently limit the solution to a singular policy for a predetermined weight, disregarding the demand for multiple trade-off solutions in real-world scenarios. When extending such approaches to find the dense Pareto set, the inefficiency of repetitively retraining the policy is unacceptable (Parisi et al. 2014). Recent studies have proposed several multi-policy methods, which simultaneously learn a set of policies over multiple preferences using a single network (Parisi et al. 2017; Chen et al. 2019; Xu et al. 2020; Basaklar et al. 2023). However, these methods either cannot construct a continuous Pareto front, or will suffer from the curse of dimensionality and poor performance when confronting with the conflicting objectives due to the reason that they only obtain a single network for all the preferences.
@@
To address these limitations, in this work, we integrate Pareto Set Learning (PSL) (Lin et al. 2022b, a) within the context of MORL, utilizing the specialty of PSL in covering the whole preference space. Considering that the solutions to MORL problems are generated by the policy networks, we propose to employ hypernetwork (Ha and Le 2017) as the PSL model to generate a continuum of Pareto-optimal policy parameters using a decomposition-based method, and thus generate the relatively personalized policy network for each preference. Since the trained hypernetwork can generate policy parameters for every possible weight without any additional cost (e.g., evaluation and fine-tuning), our proposed framework can dynamically align with user-specified preferences (Abels et al. 2019b). It is worth noting that our method is a general MORL framework, which can be equipped with any standard RL algorithm. We also perform theoretical analysis, proving the superiority of the model capacity of PSL-MORL by using Rademacher complexity (Golowich et al. 2018), and the optimality of the obtained policy network by using Banach Fixed Point Theorem (Daskalakis et al. 2018).
@@
We conduct experiments to demonstrate the effectiveness of PSL-MORL across several benchmarks, containing the Fruit Tree Navigation (FTN) task with discrete state-action spaces (Yang et al. 2019) and multiple continuous state-action control tasks MO-MuJoCo (Todorov et al. 2012; Xu et al. 2020). Since PSL-MORL is a general framework, we employ Double Deep Q-Network (DDQN) (Van Hasselt et al. 2016) for discrete environments and Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto et al. 2018b) for continuous environments as the lower-level RL algorithms. The experimental results demonstrate that our method is superior to state-of-the-art MORL methods, and the ablation study also verifies the effectiveness of the parameter fusion technique employed by PSL-MORL.
@@
In single-objective settings, the expected return offers a complete ordering over the policy space, i.e., for any two policies $\pi$ and $\pi^{\prime}$, $J^{\pi}$ will either be greater than, equal to, or lower than $J^{\pi^{\prime}}$. Thus, it is sufficient to find an optimal policy $\pi^{*}$ that maximizes the expected cumulative discounted reward. However, in MORL, there exists no single optimal policy that can maximize all the objectives simultaneously, and we need the following Pareto concepts (Qian et al. 2013, 2019; Zhou et al. 2019).
@@
Previous MORL approaches can be mainly divided into single-policy, multi-policy, and meta-policy approaches. Single-policy approaches transform a multi-objective problem into a single-objective problem by combining rewards into a single scalar reward using a utility function, and then use standard RL methods to maximize the scalarized return (Roijers et al. 2013). However, these approaches require domain-specific knowledge and predefined preferences (Van Moffaert et al. 2013; Abdolmaleki et al. 2020). Multi-policy approaches aim to obtain a good approximated Pareto set, and the most widely used approach is running a single-policy algorithm repeatedly over various preferences (Roijers et al. 2014a; Mossalam et al. 2016; Zuluaga et al. 2016). A more recent study proposes an efficient evolutionary learning algorithm to update a population of policies simultaneously in each run to improve the approximation to the Pareto front (Xu et al. 2020). However, it suffers from the low efficiency of acquiring dense Pareto optimal solutions and handling a large number of objectives (Pirotta et al. 2015; Parisi et al. 2017; Chen et al. 2019).
@@
Recent studies have proposed multi-objective Q-learning approaches that simultaneously learn a set of policies over multiple preferences, using a single network that takes preferences as inputs and uses vectorized value function updates (Abels et al. 2019b; Yang et al. 2019). Another orthogonal approach frames MORL as a meta-learning problem using a task distribution given by a distribution over the preference space (Chen et al. 2019). It first trains a meta-policy to approximate the Pareto front implicitly, and then obtains the Pareto optimal solution of a given preference by fine-tuning the meta-policy with a few gradient updates. But it often fails because the meta-policy may not be optimal, leading to the ultimate policy being suboptimal. PD-MORL (Basaklar et al. 2023) is a recently proposed method that utilizes a universal neural network to obtain the entire Pareto front. However, these approaches obtain a single policy network for all the trade-off weights, making it challenging to simultaneously address all scalarized subproblems. Compared to previous methods, our proposed PSL-MORL method generates a policy network for each preference, thereby alleviating conflicts and enhancing the overall quality of the solution set.
@@
Multi-Objective Evolutionary Algorithms (MOEAs), inspired by natural evolution, are another sort of algorithms that can solve MORL problems. They are competitive approaches in solving complicated multi-objective optimization problems (Qian et al. 2015a, b, c, 2017; Hong et al. 2019) and can be applied to some continuous control problems (Qian and Yu 2021). They obtain a high quality solution set by iteratively using crossover, mutation, and selection operators. For instance, MOEA/D (MOEA based on Decomposition) (Zhang and Li 2007) is one of the most common MOEA algorithms, which partitions a multi-objective optimization problem into a series of single-objective optimization or less complex multi-objective optimization components. It then employs a heuristic search approach to concurrently and collaboratively refine these easier optimization sub-tasks. However, MOEAs face the challenges in scaling up with the dimensionality of parameters and are not ideally suited for the training of large-scale neural networks in RL.
@@
which can be computed using stochastic optimization, i.e., $\min_{\theta\in\Theta}\sum_{i}g_{WS}(x=h_{\theta}(\lambda_{i})\mid\lambda_{i}),\{\lambda_{i}\}\sim\Lambda$. PSL has been successfully applied to multi-objective Bayesian optimization (Lin et al. 2022b; Lu et al. 2024), multi-objective neural combinatorial optimization (Lin et al. 2022a), and drug design (Jain et al. 2023). Recently, (Zhang et al. 2023) provided a novel geometric perspective for PSL and recognized its equivalence to hypervolume maximization.
@@
In this section, we introduce the Pareto Set Learning method for MORL (PSL-MORL), an efficient framework for approximating the whole Pareto front of MORL. The overall workflow of PSL-MORL is intuitively illustrated in Figure 1. PSL-MORL consists of two primary modules: a hypernetwork $\phi$ and a main policy network. The hypernetwork, when given a preference vector $\bm{\omega}$ as input, generates the parameters for the main policy network. This policy network is then used to control an agent in response to various states. During each training episode, a random preference vector $\bm{\omega}$ is drawn from a uniform distribution. After inputting the preference into the hypernetwork, the agent is run in the environment using the generated policy network parameters $\phi(\omega)$. The hypernetwork parameters, along with the target network, are subsequently updated through gradient backpropagation of the RL loss. In practice, to enhance training stability, we co-train another policy network with the hypernetwork using the parameter fusion technique (Ortiz et al. 2024).
In the following, we will first introduce the details of the general PSL-MORL framework. Then, we give the instantiations of PSL-MORL with DDQN, and prove the optimality of the generated policy network. Finally, we theoretically analyze the model capacity of PSL-MORL, and prove its superiority over the state-of-the-art method PD-MORL (Basaklar et al. 2023), which thus provides theoretical justification for its greater ability of generating more personalized policies for various preference vectors.
@@
As observed in (Ortiz et al. 2024), the performance of hypernetwork prediction has large fluctuations and is highly sensitive to the norm of the input vector, thus leading to unstable training. To address this issue and stabilize the hypernetwork training, we design a parameter fusion technique to blend the parameters predicted by the hypernetwork and the main policy network parameters, i.e.,
@@
In this subsection, we introduce an instantiation of PSL-MORL by taking DDQN (Van Hasselt et al. 2016) as the single-objective RL algorithm. The pseudo code is shown in Appendix A.1 due to space limitation. The chosen replay buffer is the Hindsight Experience Replay (HER) (Andrychowicz et al. 2017), which is used to overcome the bias towards over-represented preferences. Inspired by (Basaklar et al. 2023), we adopt the idea of dividing the preference space into several subspaces to improve the sample efficiency of the algorithm and using the multi-dimensional preference interpolator $I(\bm{\omega})$ to align the generated solution with the preference vector.
PSL-MORL with DDQN first initializes the hypernetwork and HER. The parallel child processes are the same as that in Algorithm 1. After that, it operates on a batch of preferences. For each preference, the multi-dimensional interpolator $I(\bm{\omega})$ is used to obtain the interpolated preference $\bm{w}_{p}$, and a modified Q-learning update scheme (with the cosine similarity term $S_{c}$ proposed in (Basaklar et al. 2023)) is used to calculate the target value. The empirical expected loss is then obtained by calculating the DDQN loss for each preference. Finally, an optimizer (e.g., SGD) is employed to update the parameters of both the hypernetwork and target network.
The following theorem (Yang et al. 2019; Basaklar et al. 2023) shows that, the $\mathbf{Q}$-network space of the policy is a complete metric space, and the existence of a contraction mapping $\mathcal{C}$ ensures that $\mathbf{Q}^{*}$ under the optimal policy of any preference is the unique fixed point of this contraction mapping. In other words, for PSL-MORL using DDQN as the basis RL algorithm, the generated policy network can converge to the optimal policy under any preference $\bm{\omega}$.
@@
(Yang et al. 2019; Basaklar et al. 2023) $(\mathcal{Q},d)$ is a complete metric space, where
@@
In the experiments, we also adopt PSL-MORL with TD3 (Dankwa and Zheng 2019), whose details are shown in Appendix A.2. Here, we use the hypernetwork to generate only the parameters of actor network and also conduct the parameter fusion technique. Similar to (Basaklar et al. 2023), we include the directional angle term to both the actor and critic loss function, which depicts the alignment relationship between the preference vectors and the Q-values.
@@
In this subsection, we theoretically analyze the model representation ability of PSL-MORL, which is characterized by the Rademacher complexity in Definition 4. A model with higher Rademacher complexity will have a larger model capacity (Vapnik 2000; Zhu et al. 2009; Shalev-Shwartz and Ben-David 2014; Neyshabur et al. 2019).
@@
(Rademacher Complexity (Golowich et al. 2018)) Given a real-valued function class $\mathcal{H}$ and a set of data points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\in\mathcal{X}$, the (empirical) Rademacher complexity ^$\hat{\mathcal{R}}_{n}(\mathcal{H})$ is defined as
@@
We prove in Theorem 2 that PSL-MORL has a higher Rademacher complexity (thus a better model capacity) than PD-MORL (Basaklar et al. 2023), the most competitive baseline among all the MORL algorithms. This implies a greater ability of PSL-MORL to provide more personalized policies for various preference vectors, as will also be shown in the experiments. The theoretical analysis relies on the common assumption that the activation function $\sigma$ of a neural network satisfies the 1-Lipschitz property and is positive homogeneous.
@@
To examine the performance of PSL-MORL, we conduct experiments on two popular MORL benchmarks: a continuous benchmark MO-MuJoCo (Xu et al. 2020) and a discrete benchmark Fruit Tree Navigation (FTN) (Yang et al. 2019). MO-MuJoCo is a popular MORL benchmark based on the MoJocCo physics simulation environment, consisting of several continuous control tasks. We conduct experiments under five different environments, including MO-HalfCheetah-v2, MO-Hopper-v2, MO-Ant-v2, MO-Swimmer-v2, and MO-Walker-v2. The number of objectives is two for all environments in MO-MuJoCo. FTN is a discrete MORL benchmark with six objectives, whose goal is to navigate the tree to harvest fruit to optimize six nutritional values on specific preferences. Full details of the benchmarks are provided in Appendix A.4.
To evaluate the performance of PSL-MORL, we compare the following seven state-of-the-art methods: PG-MORL (Xu et al. 2020), PD-MORL (Basaklar et al. 2023), Radial Algorithm (RA) (Parisi et al. 2014), Pareto Front Adaption (PFA) (Parisi et al. 2014), MOEA/D (Zhang and Li 2007), RANDOM (Xu et al. 2020), and META (Chen et al. 2019). Please refer to Appendix A.5 for details.
An MORL algorithm aims to find a set of policies to approximate the underlying optimal Pareto set and Pareto front. Following previous works (Chen et al. 2019; Xu et al. 2020; Liu and Qian 2021; Basaklar et al. 2023), we use two metrics: (i) Hypervolume and (ii) Sparsity (Hayes et al. 2022), to compare the quality of the set of policies obtained by different methods.
@@
The hypervolume metric can measure the convergence performance of the approximated Pareto front. However, it cannot fully describe the density of the approximated Pareto front. A dense policy set is always favored for better Pareto approximation (Chen et al. 2019; Xu et al. 2020; Liu and Qian 2021; Basaklar et al. 2023). Therefore, it is necessary to consider the sparsity metric, which measures the density of the approximated Pareto front in a multi-objective control problem. As a larger hypervolume implies a better approximated Pareto set and a lower sparsity indicates a denser approximated Pareto set, the goal is to obtain an approximated Pareto set with a high hypervolume metric and meanwhile a low sparsity metric. Note that hypervolume and sparsity are the most frequently used metrics to measure the performance of MORL algorithms (Chen et al. 2019; Xu et al. 2020; Liu and Qian 2021; Basaklar et al. 2023).
@@
We first evaluate the methods on the popular multi-objective continuous control benchmark MO-MuJoCo. We run each method with six different random seeds, consistent with previous works (Xu et al. 2020; Basaklar et al. 2023). We do not compare with Envelope Q-Learning (Yang et al. 2019) on MO-MuJoCo since it cannot be applied to continuous action spaces. For a fair comparison, we use the same seeds and hyperparameters for other baseline algorithms. As shown in Table 1, PSL-MORL achieves the best average rank for both metrics, demonstrating the high quality of the obtained Pareto set. PD-MORL is the runner-up, performing worse than PSL-MORL which may be due to the limited capacity and diversity of its single policy model. PG-MORL involves a large optimization space for training multiple policies simultaneously, achieving the third average rank. RANDOM performs worse than PG-MORL, because it does not consider the influence of the choice of weights. Due to the fixed weights and the low efficiency of training separately, RA doesn’t have a good performance on the hypervolume metric. PFA also performs poorly since it trains the policy separately and the transfer may not be optimal. The low sample efficiency of the gradient-free optimization in MOEA/D leads to its poor performance. META has the worst hypervolume, because it can only obtain a sub-optimal policy.
@@
We evaluate the proposed PSL-MORL on the commonly used benchmark Fruit Tree Navigation (FTN) with different depths, i.e., $d=5,6,7$. As shown in Table 3, PSL-MORL achieves the best performance in both metrics in all the environments. When the depths are $5$ and $6$, Envelope (Yang et al. 2019) and PD-MORL (Basaklar et al. 2023) fail to find the whole Pareto set but find only one policy, resulting in an N/A sparsity value. However, PSL-MORL can find a better approximated Pareto set and get a $0.01$ sparsity. The results demonstrate that PSL-MORL has a better exploration capability than Envelope and PD-MORL. Furthermore, as the depth increases from $5$ to $7$, the hypervolume gap between PSL-MORL and the other two methods also increases, highlighting the greater advantage in more complex environments.

### v2

@@ Introduction
Reinforcement Learning (RL) (Sutton and Barto 2018) has become a prevalent technique for addressing decision-making and control problems in a wide range of scenarios, including video games (Mnih et al. 2015), navigation (Yang, Sun, and Narasimhan 2019), robotics (Schulman et al. 2017), dynamic algorithm configuration (Xue et al. 2022), and financial cloud services (Yang et al. 2024). In standard RL, the main objective is obtaining an optimal policy that achieves the best expected returns. However, in many real-world scenarios, spanning from robotics, where the balance between speed and energy efficiency is paramount (Todorov, Erez, and Tassa 2012), to smart systems design, where the trade-offs between competing metrics must be dynamically managed (Azzouz et al. 2018), the problems may involve multiple conflicting objectives. Contrary to standard RL, Multi-Objective RL (MORL) aims to generate an optimal policy tailored to each preference over multiple objectives, constructing a diverse set of optimal policies that collectively form a Pareto set.
@@
Traditional MORL methods transform the multi-dimensional objective space into a scalar representation (Mannor and Shimkin 2001) through the static assignment of weights to each objective, achieving a trade-off between multiple objectives. Such approaches, however, struggle with objectives of varying magnitudes and require extensive domain knowledge to set the appropriate weights/preferences (Amodei et al. 2016). Moreover, they inherently limit the solution to a singular policy for a predetermined weight, disregarding the demand for multiple trade-off solutions in real-world scenarios. When extending such approaches to find the dense Pareto set, the inefficiency of repetitively retraining the policy is unacceptable (Parisi et al. 2014). Recent studies have proposed several multi-policy methods, which simultaneously learn a set of policies over multiple preferences using a single network (Parisi, Pirotta, and Peters 2017; Chen et al. 2019; Xu et al. 2020; Basaklar, Gumussoy, and Ogras 2023). However, these methods either cannot construct a continuous Pareto front, or will suffer from the curse of dimensionality and poor performance when confronting with the conflicting objectives due to the reason that they only obtain a single network for all the preferences.
@@
To address these limitations, in this work, we integrate Pareto Set Learning (PSL) (Lin et al. 2022; Lin, Yang, and Zhang 2022) within the context of MORL, utilizing the specialty of PSL in covering the whole preference space. Considering that the solutions to MORL problems are generated by the policy networks, we propose to employ hypernetwork (Ha and Le 2017) as the PSL model to generate a continuum of Pareto-optimal policy parameters using a decomposition-based method, and thus generate the relatively personalized policy network for each preference. Since the trained hypernetwork can generate policy parameters for every possible weight without any additional cost (e.g., evaluation and fine-tuning), our proposed framework can dynamically align with user-specified preferences (Abels et al. 2019). It is worth noting that our method is a general MORL framework, which can be equipped with any standard RL algorithm. We also perform theoretical analysis, proving the superiority of the model capacity of PSL-MORL by using Rademacher complexity (Golowich, Rakhlin, and Shamir 2018), and the optimality of the obtained policy network by using Banach Fixed Point Theorem (Daskalakis, Tzamos, and Zampetakis 2018).
@@
We conduct experiments to demonstrate the effectiveness of PSL-MORL across several benchmarks, containing the Fruit Tree Navigation (FTN) task with discrete state-action spaces (Yang, Sun, and Narasimhan 2019) and multiple continuous state-action control tasks MO-MuJoCo (Todorov, Erez, and Tassa 2012; Xu et al. 2020). Since PSL-MORL is a general framework, we employ Double Deep Q-Network (DDQN) (Van Hasselt, Guez, and Silver 2016) for discrete environments and Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto, Hoof, and Meger 2018) for continuous environments as the lower-level RL algorithms. The experimental results demonstrate that our method is superior to state-of-the-art MORL methods, and the ablation study also verifies the effectiveness of the parameter fusion technique employed by PSL-MORL.
@@
In single-objective settings, the expected return offers a complete ordering over the policy space, i.e., for any two policies $\pi$ and $\pi^{\prime}$, $J^{\pi}$ will either be greater than, equal to, or lower than $J^{\pi^{\prime}}$. Thus, it is sufficient to find an optimal policy $\pi^{*}$ that maximizes the expected cumulative discounted reward. However, in MORL, there exists no single optimal policy that can maximize all the objectives simultaneously, and we need the following Pareto concepts (Qian, Yu, and Zhou 2013; Qian et al. 2019; Zhou, Yu, and Qian 2019).
@@
Previous MORL approaches can be mainly divided into single-policy, multi-policy, and meta-policy approaches. Single-policy approaches transform a multi-objective problem into a single-objective problem by combining rewards into a single scalar reward using a utility function, and then use standard RL methods to maximize the scalarized return (Roijers et al. 2013). However, these approaches require domain-specific knowledge and predefined preferences (Van Moffaert, Drugan, and Nowé 2013; Abdolmaleki et al. 2020). Multi-policy approaches aim to obtain a good approximated Pareto set, and the most widely used approach is running a single-policy algorithm repeatedly over various preferences (Roijers et al. 2014; Mossalam et al. 2016; Zuluaga, Krause, and Püschel 2016). A more recent study proposes an efficient evolutionary learning algorithm to update a population of policies simultaneously in each run to improve the approximation to the Pareto front (Xu et al. 2020). However, it suffers from the low efficiency of acquiring dense Pareto optimal solutions and handling a large number of objectives (Pirotta, Parisi, and Restelli 2015; Parisi, Pirotta, and Peters 2017; Chen et al. 2019).
@@
Recent studies have proposed multi-objective Q-learning approaches that simultaneously learn a set of policies over multiple preferences, using a single network that takes preferences as inputs and uses vectorized value function updates (Abels et al. 2019; Yang, Sun, and Narasimhan 2019). Another orthogonal approach frames MORL as a meta-learning problem using a task distribution given by a distribution over the preference space (Chen et al. 2019). It first trains a meta-policy to approximate the Pareto front implicitly, and then obtains the Pareto optimal solution of a given preference by fine-tuning the meta-policy with a few gradient updates. But it often fails because the meta-policy may not be optimal, leading to the ultimate policy being suboptimal. PD-MORL (Basaklar, Gumussoy, and Ogras 2023) is a recently proposed method that utilizes a universal neural network to obtain the entire Pareto front. However, these approaches obtain a single policy network for all the trade-off weights, making it challenging to simultaneously address all scalarized subproblems. Compared to previous methods, our proposed PSL-MORL method generates a policy network for each preference, thereby alleviating conflicts and enhancing the overall quality of the solution set.
@@
Multi-Objective Evolutionary Algorithms (MOEAs), inspired by natural evolution, are another sort of algorithms that can solve MORL problems. They are competitive approaches in solving complicated multi-objective optimization problems (Qian, Yu, and Zhou 2015a, b, c; Qian et al. 2017; Hong et al. 2019) and can be applied to some continuous control problems (Qian and Yu 2021). They obtain a high quality solution set by iteratively using crossover, mutation, and selection operators. For instance, MOEA/D (MOEA based on Decomposition) (Zhang and Li 2007) is one of the most common MOEA algorithms, which partitions a multi-objective optimization problem into a series of single-objective optimization or less complex multi-objective optimization components. It then employs a heuristic search approach to concurrently and collaboratively refine these easier optimization sub-tasks. However, MOEAs face the challenges in scaling up with the dimensionality of parameters and are not ideally suited for the training of large-scale neural networks in RL.
@@
which can be computed using stochastic optimization, i.e., $\min_{\theta\in\Theta}\sum_{i}g_{WS}(x=h_{\theta}(\lambda_{i})\mid\lambda_{i}),\{\lambda_{i}\}\sim\Lambda$. PSL has been successfully applied to multi-objective Bayesian optimization (Lin et al. 2022; Lu, Li, and Zhou 2024), multi-objective neural combinatorial optimization (Lin, Yang, and Zhang 2022), and drug design (Jain et al. 2023). Recently, (Zhang et al. 2023) provided a novel geometric perspective for PSL and recognized its equivalence to hypervolume maximization.
@@
In this section, we introduce the Pareto Set Learning method for MORL (PSL-MORL), an efficient framework for approximating the whole Pareto front of MORL. The overall workflow of PSL-MORL is intuitively illustrated in Figure 1. PSL-MORL consists of two primary modules: a hypernetwork $\phi$ and a main policy network. The hypernetwork, when given a preference vector $\bm{\omega}$ as input, generates the parameters for the main policy network. This policy network is then used to control an agent in response to various states. During each training episode, a random preference vector $\bm{\omega}$ is drawn from a uniform distribution. After inputting the preference into the hypernetwork, the agent is run in the environment using the generated policy network parameters $\phi(\omega)$. The hypernetwork parameters, along with the target network, are subsequently updated through gradient backpropagation of the RL loss. In practice, to enhance training stability, we co-train another policy network with the hypernetwork using the parameter fusion technique (Ortiz, Guttag, and Dalca 2024).
In the following, we will first introduce the details of the general PSL-MORL framework. Then, we give the instantiations of PSL-MORL with DDQN, and prove the optimality of the generated policy network. Finally, we theoretically analyze the model capacity of PSL-MORL, and prove its superiority over the state-of-the-art method PD-MORL (Basaklar, Gumussoy, and Ogras 2023), which thus provides theoretical justification for its greater ability of generating more personalized policies for various preference vectors.
@@
As observed in (Ortiz, Guttag, and Dalca 2024), the performance of hypernetwork prediction has large fluctuations and is highly sensitive to the norm of the input vector, thus leading to unstable training. To address this issue and stabilize the hypernetwork training, we design a parameter fusion technique to blend the parameters predicted by the hypernetwork and the main policy network parameters, i.e.,
@@
In this subsection, we introduce an instantiation of PSL-MORL by taking DDQN (Van Hasselt, Guez, and Silver 2016) as the single-objective RL algorithm. The pseudo code is shown in Appendix A.1 due to space limitation. The chosen replay buffer is the Hindsight Experience Replay (HER) (Andrychowicz et al. 2017), which is used to overcome the bias towards over-represented preferences. Inspired by (Basaklar, Gumussoy, and Ogras 2023), we adopt the idea of dividing the preference space into several subspaces to improve the sample efficiency of the algorithm and using the multi-dimensional preference interpolator $I(\bm{\omega})$ to align the generated solution with the preference vector.
PSL-MORL with DDQN first initializes the hypernetwork and HER. The parallel child processes are the same as that in Algorithm 1. After that, it operates on a batch of preferences. For each preference, the multi-dimensional interpolator $I(\bm{\omega})$ is used to obtain the interpolated preference $\bm{w}_{p}$, and a modified Q-learning update scheme (with the cosine similarity term $S_{c}$ proposed in (Basaklar, Gumussoy, and Ogras 2023)) is used to calculate the target value. The empirical expected loss is then obtained by calculating the DDQN loss for each preference. Finally, an optimizer (e.g., SGD) is employed to update the parameters of both the hypernetwork and target network.
The following theorem (Yang, Sun, and Narasimhan 2019; Basaklar, Gumussoy, and Ogras 2023) shows that, the $\mathbf{Q}$-network space of the policy is a complete metric space, and the existence of a contraction mapping $\mathcal{C}$ ensures that $\mathbf{Q}^{*}$ under the optimal policy of any preference is the unique fixed point of this contraction mapping. In other words, for PSL-MORL using DDQN as the basis RL algorithm, the generated policy network can converge to the optimal policy under any preference $\bm{\omega}$.
@@
(Yang, Sun, and Narasimhan 2019; Basaklar, Gumussoy, and Ogras 2023) $(\mathcal{Q},d)$ is a complete metric space, where
@@
In the experiments, we also adopt PSL-MORL with TD3 (Dankwa and Zheng 2019), whose details are shown in Appendix A.2. Here, we use the hypernetwork to generate only the parameters of actor network and also conduct the parameter fusion technique. Similar to (Basaklar, Gumussoy, and Ogras 2023), we include the directional angle term to both the actor and critic loss function, which depicts the alignment relationship between the preference vectors and the Q-values.
@@
In this subsection, we theoretically analyze the model representation ability of PSL-MORL, which is characterized by the Rademacher complexity in Definition 4. A model with higher Rademacher complexity will have a larger model capacity (Vapnik 2000; Zhu, Rogers, and Gibson 2009; Shalev-Shwartz and Ben-David 2014; Neyshabur et al. 2019).
@@
(Rademacher Complexity (Golowich, Rakhlin, and Shamir 2018)) Given a real-valued function class $\mathcal{H}$ and a set of data points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\in\mathcal{X}$, the (empirical) Rademacher complexity ^$\hat{\mathcal{R}}_{n}(\mathcal{H})$ is defined as
@@
We prove in Theorem 2 that PSL-MORL has a higher Rademacher complexity (thus a better model capacity) than PD-MORL (Basaklar, Gumussoy, and Ogras 2023), the most competitive baseline among all the MORL algorithms. This implies a greater ability of PSL-MORL to provide more personalized policies for various preference vectors, as will also be shown in the experiments. The theoretical analysis relies on the common assumption that the activation function $\sigma$ of a neural network satisfies the 1-Lipschitz property and is positive homogeneous.
@@
To examine the performance of PSL-MORL, we conduct experiments on two popular MORL benchmarks: a continuous benchmark MO-MuJoCo (Xu et al. 2020) and a discrete benchmark Fruit Tree Navigation (FTN) (Yang, Sun, and Narasimhan 2019). MO-MuJoCo is a popular MORL benchmark based on the MoJocCo physics simulation environment, consisting of several continuous control tasks. We conduct experiments under five different environments, including MO-HalfCheetah-v2, MO-Hopper-v2, MO-Ant-v2, MO-Swimmer-v2, and MO-Walker-v2. The number of objectives is two for all environments in MO-MuJoCo. FTN is a discrete MORL benchmark with six objectives, whose goal is to navigate the tree to harvest fruit to optimize six nutritional values on specific preferences. Full details of the benchmarks are provided in Appendix A.4.
To evaluate the performance of PSL-MORL, we compare the following seven state-of-the-art methods: PG-MORL (Xu et al. 2020), PD-MORL (Basaklar, Gumussoy, and Ogras 2023), Radial Algorithm (RA) (Parisi et al. 2014), Pareto Front Adaption (PFA) (Parisi et al. 2014), MOEA/D (Zhang and Li 2007), RANDOM (Xu et al. 2020), and META (Chen et al. 2019). Please refer to Appendix A.5 for details.
An MORL algorithm aims to find a set of policies to approximate the underlying optimal Pareto set and Pareto front. Following previous works (Chen et al. 2019; Xu et al. 2020; Liu and Qian 2021; Basaklar, Gumussoy, and Ogras 2023), we use two metrics: (i) Hypervolume and (ii) Sparsity (Hayes et al. 2022), to compare the quality of the set of policies obtained by different methods.
@@
The hypervolume metric can measure the convergence performance of the approximated Pareto front. However, it cannot fully describe the density of the approximated Pareto front. A dense policy set is always favored for better Pareto approximation (Chen et al. 2019; Xu et al. 2020; Liu and Qian 2021; Basaklar, Gumussoy, and Ogras 2023). Therefore, it is necessary to consider the sparsity metric, which measures the density of the approximated Pareto front in a multi-objective control problem. As a larger hypervolume implies a better approximated Pareto set and a lower sparsity indicates a denser approximated Pareto set, the goal is to obtain an approximated Pareto set with a high hypervolume metric and meanwhile a low sparsity metric. Note that hypervolume and sparsity are the most frequently used metrics to measure the performance of MORL algorithms (Chen et al. 2019; Xu et al. 2020; Liu and Qian 2021; Basaklar, Gumussoy, and Ogras 2023).
@@
We first evaluate the methods on the popular multi-objective continuous control benchmark MO-MuJoCo. We run each method with six different random seeds, consistent with previous works (Xu et al. 2020; Basaklar, Gumussoy, and Ogras 2023). We do not compare with Envelope Q-Learning (Yang, Sun, and Narasimhan 2019) on MO-MuJoCo since it cannot be applied to continuous action spaces. For a fair comparison, we use the same seeds and hyperparameters for other baseline algorithms. As shown in Table 1, PSL-MORL achieves the best average rank for both metrics, demonstrating the high quality of the obtained Pareto set. PD-MORL is the runner-up, performing worse than PSL-MORL which may be due to the limited capacity and diversity of its single policy model. PG-MORL involves a large optimization space for training multiple policies simultaneously, achieving the third average rank. RANDOM performs worse than PG-MORL, because it does not consider the influence of the choice of weights. Due to the fixed weights and the low efficiency of training separately, RA doesn’t have a good performance on the hypervolume metric. PFA also performs poorly since it trains the policy separately and the transfer may not be optimal. The low sample efficiency of the gradient-free optimization in MOEA/D leads to its poor performance. META has the worst hypervolume, because it can only obtain a sub-optimal policy.
@@
We evaluate the proposed PSL-MORL on the commonly used benchmark Fruit Tree Navigation (FTN) with different depths, i.e., $d=5,6,7$. As shown in Table 3, PSL-MORL achieves the best performance in both metrics in all the environments. When the depths are $5$ and $6$, Envelope (Yang, Sun, and Narasimhan 2019) and PD-MORL (Basaklar, Gumussoy, and Ogras 2023) fail to find the whole Pareto set but find only one policy, resulting in an N/A sparsity value. However, PSL-MORL can find a better approximated Pareto set and get a $0.01$ sparsity. The results demonstrate that PSL-MORL has a better exploration capability than Envelope and PD-MORL. Furthermore, as the depth increases from $5$ to $7$, the hypervolume gap between PSL-MORL and the other two methods also increases, highlighting the greater advantage in more complex environments.